OpenStudy (anonymous):

Suppose that the number of bacteria in a certain population increases according to an exponential growth model. A sample of 2600 bacteria selected from this population reached the size of 2802 bacteria in three hours. Find the continuous growth rate per hour. Write your answer as a percentage. Do not round any intermediate computations, and round your percentage to the nearest hundredth.

OpenStudy (anonymous):

exponential growth is of the form a*b^t here a is the initial amount 2600. t is the time in hours. We know 2600*b^3=2802 from the given information. We solve for the growth rate b: (2802/2600)^(1/3)=b=1.0253 so we are growing at .0253*100=2.53% per hour